# Localizing Factual Inconsistencies in Attributable Text Generation

### v1

@@ 8 Conclusion
We introduced QASemConsistency, a methodology that detects and localizes factual consistency errors to specific predicate-argument relations. Our error localization method is robust, as shown by our high human agreement. Contrast with other methods that either created complex, non-granular claims in natural language, or relied on linguistic formalisms that excluded non-expert annotators, our method can be applied with ease by non-expert users and models alike, without sacrificing granularity. Moreover, it can help human consumers of generative models recognize which parts of a response should be double checked. We hope that future work will focus on extending our approach to detect inconsistencies in the wider discourse, and help generative models to correct their responses.
@@ Acknowledgments
We thank all the Mturk workers that participate in our annotation project. We thank Eran Hirsch, Aviv Slobodkin, Ayal Klein, Shuyang Cao, Vidhisha Balachandran, Tanya Goyal, Greg Durrett, Kyle Lo, Philippe Laban for their valuable discussions at various stages of the project. This research has been supported by a Google gift grant and the Israel Science Foundation (grant no. 2827/21). We also thank Google Cloud for providing us with credits for running experiments on the Google Cloud Platform. Arie Cattan is partially supported by the PBC fellowship for outstanding PhD candidates in data science.
Appendix A QASem parser
To automatically generate the QASem QAs for each generated text, we re-train the joint QASem parser Klein et al. (2022) and replace the original T5-small model with T5-XL (3B). This parser takes as input a sentence and a predicate, whether verbal or nominal, and generates a list of atomic QAs. We train our parser for 5 epochs until convergence with Adam optimizer and a learning rate of $055e-$ 05. We evaluate our parser on the QASRL gold data (Roit et al., 2020) and QANom (Klein et al., 2020), achieving 75.9 F1 (+7.3) on QASRL and 72.4 F1 (+13.2) on QANom. Similarly to (Klein et al., 2022), we use a span match threshold of IOU >= 0.3 to match between predicted and gold arguments.
Appendix B Human Evaluation
@@
Although our QASem parser significantly outperforms the current SOTA model (Klein et al., 2022) by a large margin (see §A), it is still prone to errors that can introduce noise into the consistency evaluation. Specifically, for nominal predicates, the parser may generate QAs that are nonsensical or convey a different meaning from the intended predicate. For instance, in the sentence “Sweden’s former foreign minister Johan Gustafssoon, who was kidnapped by Islamist militants in Mali in 2013, has been released after more than two years in captivity” (from CLIFF) the parser incorrectly generates ‘Where did someone militantise?” for the predicate “militants” and ‘Who was captained?” for the predicate “captivity”. In some rare cases, the parser hallucinates predicate argument relation. For instance, in the sentence “A man has been arrested on suspicion of murder after a man was found dead at a house in Cambridgeshire”, the parser generates the question “Where was someone arrested? at a house in Cambridgeshire”, while this location refers to the predicate “found”. To mitigate this issue, we add a preliminary annotation step, in which an annotator was asked to verify whether the QAs are both (1) semantically interpretable and (2) correctly represent the semantic relations in the generated text. This preliminary step is efficient and required less than an hour to validate 500 QAs.

### v2

@@ 8 Conclusion
We introduced QASemConsistency, a methodology that detects and localizes factual consistency errors to specific predicate-argument relations. Our error localization method is robust, as shown by our high human agreement and strong correlation with human preferences. Contrast with other methods that either created complex, non-granular claims in natural language, or relied on linguistic formalisms that excluded non-expert annotators, our method can be applied with ease by non-expert users and models alike. Moreover, it can help human consumers of generative models recognize which parts of a response should be double checked.
We hope that future work will focus on extending our approach to detect inconsistencies in the wider discourse, and help generative models to correct their responses.
@@ Acknowledgments
We thank the anonymous reviewers and the action editor for their valuable suggestions. We also thank all the Mturk workers that participate in our annotation project. Special thanks to Eran Hirsch, Aviv Slobodkin, Ayal Klein, Shuyang Cao, Vidhisha Balachandran, Tanya Goyal, Greg Durrett, Kyle Lo, Philippe Laban for their valuable discussions at various stages of the project and Ron Eliav for his considerable help in training the QASem parser. This research has been supported by a Google gift grant and the Israel Science Foundation (grant no. 2827/21). We also thank Google Cloud for providing us with credits for running experiments on the Google Cloud Platform. Arie Cattan is partially supported by the PBC fellowship for outstanding PhD candidates in data science.
Appendix A Dataset Collection
@@
Although our QASem parser significantly outperforms the current SOTA model (Klein et al., 2022) by a large margin (see §3), it is still prone to errors that can introduce noise into the consistency evaluation. Specifically, for nominal predicates, the parser may generate QAs that are nonsensical or convey a different meaning from the intended predicate. For instance, in the sentence “Sweden’s former foreign minister Johan Gustafssoon, who was kidnapped by Islamist militants in Mali in 2013, has been released after more than two years in captivity” (from CLIFF) the parser incorrectly generates ‘Where did someone militantise?” for the predicate “militants” and ‘Who was captained?” for the predicate “captivity”. In some rare cases, the parser hallucinates predicate argument relation. For instance, in the sentence “A man has been arrested on suspicion of murder after a man was found dead at a house in Cambridgeshire”, the parser generates the question “Where was someone arrested? at a house in Cambridgeshire”, while this location refers to the predicate “found”. To mitigate this issue, we add a preliminary annotation step, in which an annotator was asked to verify whether the QAs are both (1) semantically interpretable and (2) correctly represent the semantic relations in the generated text. This preliminary step is efficient and required less than an hour to validate 500 QAs.
@@
Annotator Compensation
For XSUM and biographies, each HIT consists of one or reference text, $\mathcal{X}$, and two model outputs, $\mathcal{Y}_{1}$ and $\mathcal{Y}_{2}$. For XSUM, each HIT includes a source article with summaries generated by BART and PEGASUS, and compensated with \$1.50 per HIT. For biographies, we select two model outputs for each instance, with a compensation of \$3.00 per HIT, reflecting the larger number of QAs to annotate. For Verifiability, each HIT consists of a single response with a compensation of \$2.50 per HIT.
Appendix B Prompts